# MotionStream: Real-Time Video Generation with Interactive Motion Controls

### v1

@@ 4.1 Implementation Details
We build upon image to video (I2V) of variants of Wan 2.1 (1.3B) (Wang et al., 2025a) and Wan 2.2 (5B) (Wan et al., 2025). We train teacher models on OpenVid-1M (Nan et al., 2024) and synthetic data generated by Wan text-to-video (T2V) model: Wan 2.1 at 832×480 (70K synthetic samples) and Wan 2.2 at 1280×704 (30K synthetic samples). For causal adaptation and Self Forcing distillation, we sample input images, text prompts and 2D tracks from the synthetic datasets described above. For computational efficiency, we track all real and synthetic videos from a 50×50 uniform grid with CoTracker3 (Karaev et al., 2024). We refer readers to Appendix for additional details.
@@ 4.2 Quantitative Evaluations
Image Conductor (Li et al., 2025d) AnimateDiff (256P) 2.98 11.30 0.214 0.664 91.64 10.29 0.192 0.644 31.22
Go-With-The-Flow Burgert et al. (2025) CogVideoX-5B (480P) 0.60 15.62 0.392 0.490 41.99 14.59 0.410 0.425 10.27
Diffusion-As-Shader (Gu et al., 2025b) CogVideoX-5B (480P) 0.29 15.80 0.372 0.483 40.23 14.51 0.382 0.437 18.76
ATI (Wang et al., 2025b) Wan 2.1-14B (480P) 0.23 15.33 0.374 0.473 17.41 16.04 0.502 0.366 6.12
@@
Motion Transfer. We evaluate motion-following capability on two datasets: the DAVIS (Perazzi et al., 2016) validation set (30 videos) and 20 curated videos from the Sora webpage (Brooks et al., 2024). We use both because DAVIS presents challenging sequences with significant occlusions, while the Sora set provides clean examples with consistent visibility, ensuring a comprehensive evaluation. We directly compare the synthesized results with the corresponding ground-truth video frames. Visual fidelity is measured using PSNR, SSIM, and LPIPS (Zhang et al., 2018), while motion accuracy is assessed via End-Point Error (EPE), computed as the L2 distance between visible input tracks and the tracks extracted from the generated videos. All models are evaluated in its optimal configuration, and evaluations are performed at $832\times 480$ resolution after resizing. Speed measures are based on a single H100 GPU. We refer readers to Appendix for detailed protocols.
Camera Control. To assess our models’ ability to generate videos following camera controls, we evaluate its zero-shot performance on single-image 3D novel view synthesis. We compare it against several recent diffusion- and feed-forward-based view synthesis baselines (Zhou et al., 2025b; Yu et al., 2024; Xu et al., 2025) on the LLFF dataset (Mildenhall et al., 2019).
@@
DepthSplat (Xu et al., 2025) 576P 1.40 13.9 0.28 0.30
ViewCrafter (Yu et al., 2024) 576P 0.26 14.0 0.30 0.30
SEVA (Yu et al., 2024) 576P 0.20 14.1 0.30 0.29
@@
To adapt our 2D track-controlled model for this task, we first estimate scene geometry using a monocular depth network (Wang et al., 2025c) and compute a single scale factor to align the predicted depth with the reconstructed scene point cloud from COLMAP (Schonberger & Frahm, 2016). We then derive 2D motion trajectories by interpolating between the input and target frame using depth and camera parameters. For all methods, the input image aspect ratio is fixed at 16:9, and all synthesized frames are resized to $512\times 288$ before evaluation with PSNR, SSIM, and LPIPS metrics. As shown in Table 2, our track-conditioned video generation model outperforms other 3D novel view synthesis baselines by a large margin, even though it is not specifically designed for this task. Moreover, our causal models achieve significantly higher generation throughput compared with both the baselines and their bidirectional counterparts.
@@ 4.3 Ablation Experiments
Track Representation. We compare our sinusoidal position encoding with a learnable track head against the RGB encoding strategy using a frozen VAE, following prior work (Gu et al., 2025b), where each 2D track is assigned a unique RGB color vector and placed onto a canvas before being fed into the VAE. Table 3 shows that our method (PE-Head) outperforms RGB-VAE in both efficiency and quality. Specifically, our lightweight PE-based encoding achieves better motion alignment while being two orders of magnitude faster than the VAE-based approach. We hypothesize sinusoidal encoding preserves stronger identification signals compared to RGB encoding due to richer expressive dimensions.
@@
We additionaly show reconstruction accuracy and throughput as well as their time evolution in Table 4 and Figure 6(b,c) respectively. We denote our configuration as chunk-3, sink-1, window-1 (abbreviated c3s1w1), while sliding window approach used in Self Forcing is c3s0w6, attending to a maximum of 6 previous chunks without sink tokens, with unbounded RoPE positions that scale with temporal frame count. Although removing sink tokens (c3s0w1) yields marginal speed improvements, this modification comes at the cost of degraded long-term generation stability.
@@ Appendix A Training Efficient Tiny VAE
Tiny VAE (Wan 2.1  (Boer Bohan, 2025)) 9.84M 8×8×4 0.12 28.85 0.899 0.168
@@
Inspired by community implementations of Tiny VAE (Boer Bohan, 2025), we train a compact decoder from scratch with larger data, cleaner training pipelines, and better loss designs. While Boer Bohan (2025) trains Tiny VAE by regressing outputs to the original Wan VAE using adversarial loss from a PatchGAN discriminator and reconstruction loss with replay buffer, we extend this design by incorporating LPIPS loss with proper data scaling, and hyperparameter selections. We use a random subset of videos from OpenVid-1M and synthetic Wan videos (total 280K samples), training for 200K steps with a learning rate of $3\times 10^{-4}$, batch size of 16, and AdamW optimizer. We train at a lower resolution of $144\times 144$ and frame length of 21, but the model scales well to larger video dimensions.
@@ Appendix B VBench results and User study
We additionally evaluate MotionStream using VBench-I2V (Huang et al., 2024) and conduct user studies on 20 samples from the Sora demo subset. In Vbench-I2V, we exclude camera motion and dynamic degree metrics since these dimensions are already constrained by the input trajectory conditions rather than text prompts. VBench results show strong correlation with the underlying backbone model, favoring recent Wan-based architectures. The provision of image and trajectory conditions leads to uniformly high scores across methods, reducing discriminative power between models. Nonetheless, both our teacher and distilled models consistently achieve competitive performance across all evaluated dimensions.
@@
Image Conductor (Li et al., 2025d) 0.847 0.868 0.791 0.889 0.906 0.505 0.689
GWTF (Burgert et al., 2025) 0.957 0.974 0.933 0.944 0.981 0.620 0.675
DAS (Gu et al., 2025b) 0.972 0.987 0.953 0.958 0.988 0.634 0.695
ATI (Wang et al., 2025b) 0.981 0.988 0.948 0.947 0.980 0.629 0.707
@@ Appendix C Additional Ablation Experiments and Qualitative Results
Impact of Chunk Size and Sampling Steps. In the main paper, we primarily focused on evaluating the efficiency of diverse sparse attention patterns with attention sink size and attention window size. As shown in Table 4 and Figure 6, reducing the chunk size increases the number of autoregressive rollouts required to generate the same length of video while reducing the amount of bidirectional attention, leading to worse performance. Since chunk sizes beyond 3 induce high latency, which makes them less useful for our interactive use case, we report only speed metrics for these configurations. While we did not perform controlled training ablations for larger chunk sizes, we observe similar trends with TalkingMachines (Low & Wang, 2025), where quality generally improves with larger chunks at the cost of latency. Figure A2 (a) visualizes the latency and throughput given different chunk sizes and sampling steps. Note that since all benchmarks are conducted with a sink size and a window size of 1, the model with chunk size 7 has lower FPS than the one with chunk size 3 since its attention sequence length is longer.
@@
Streaming Demo. We show some examples of our streaming demo in Figure A5. The demo starts by accepting an input image and text prompts, which can help generate effects that are not achievable through mouse drags. Users can then choose the specific spacing/size of the track grids and start controlling objects in a scene, or move the camera. Due to its autoregressive nature, users can pause or resume the streaming generation process. Users can start/end or pause/resume the generation process with enter and space keyboard input, which is especially useful for dynamically adding static grids to specify unmoving regions or multiple moving grids to control different motions during streaming. As a video generative model, our method naturally supports drag-based image editing, generating intermediate transition frames as a bonus, while being faster than most dedicated drag-based image editing methods (Pan et al., 2023; Shi et al., 2024b; Nie et al., 2023; Shin et al., 2024; Zhao et al., 2024). To further support diverse downstream tasks, we will continuously update our front-end UI with additional features.
@@ Appendix D Training Details
Data Preprocessing. We train our models on two primary data sources: OpenVid-1M (Nan et al., 2024) and synthetic videos generated by larger Wan text-to-video models. For OpenVid-1M, we filter the dataset to 0.6M videos by requiring a minimum of 81 frames and a 16:9 aspect ratio, sampling at 16 FPS. For synthetic data, we use 70K samples for Wan 2.1 (81 frames, 480P resolution, generated by Wan 2.1 14B using text prompts from VidProm (Wang & Yang, 2024)) and 30K publicly available samples for Wan 2.2 (121 frames, 720P resolution, generated by Wan 2.2 5B from the FastVideo team (Team, 2024)).
We extract motion trajectories from all videos using CoTracker3 (Karaev et al., 2024), tracking points on a $\times50$ uniform grid. Our training follows a two-stage design: initial training on OpenVid-1M to establish general motion conditioning capability, followed by fine-tuning on cleaner synthetic data to improve trajectory adherence and reduce artifacts from noisy real-world videos. For the distillation phase, we use only synthetic samples since the distillation process converges quickly and requires less data.
@@
with timestep shifting (Wang et al., 2025a), defined as $t’(k,t)=\frac{kt/1000}{1+(k-1)(t/1000)}$, where $k=6.0$. The model is trained to predict the expected velocity fields conditioned on noisy latents, denoising time steps, text prompts $c_{t}$ and motion control embedding $c_{m}$, with timestep-dependent weighting $w_{t}$:
@@
Following CausVid (Yin et al., 2025), we adapt to few-step causal inference through ODE trajectory regression. We generate 4,000 ODE trajectories from the teacher model and train for 2,000 steps. While the transition from bidirectional to causal attention requires significant adaptation, we find variations within causal patterns (e.g., different window sizes) are similar enough to learn jointly with single model. We therefore train with diverse sparse causal attention masks, creating a unified initialization that supports flexible self-rollout configurations. We maintain a batch size of 128 with a learning rate $2\times 10^{-6}$.
Self Forcing Distillation with DMD.
Self Forcing distillation converges quickly at around 400 steps with a batch size of 64. We set learning rates to $2\times 10^{-6}$ for the generator and $4\times 10^{-7}$ for the critic (fake score function), with a 1:5 update ratio and gradient truncation as described in Sec. 3.2.
We employ AdamW optimizer (Kingma & Ba, 2015; Loshchilov & Hutter, 2019) with mixed precision (bfloat16) and PyTorch’s FSDP. Exponential moving average (EMA) is applied during teacher training and DMD distillation. Wan 2.1 variants train at $832\times 480$ resolution while Wan 2.2 trains at $1280\times 704$. With 32 A100 GPUs, training the Wan 2.1 teacher model takes roughly 3 days with causal adaptation and distillation completing in 20 hours, while Wan 2.2 requires slightly longer training times.

### v2

@@ 4.1 Implementation Details
We build upon image to video (I2V) of variants of Wan 2.1 (1.3B) (wang2025wan) and Wan 2.2 (5B) (wan2025). We train teacher models on OpenVid-1M (nan2024openvid) and synthetic data generated by Wan text-to-video (T2V) model: Wan 2.1 at 832×480 (70K synthetic samples) and Wan 2.2 at 1280×704 (30K synthetic samples). For causal adaptation and Self Forcing distillation, we sample input images, text prompts and 2D tracks from the synthetic datasets described above. For computational efficiency, we track all real and synthetic videos from a 50×50 uniform grid with CoTracker3 (karaev2024cotracker3). We refer readers to Appendix for additional details.
@@ 4.2 Quantitative Evaluations
Image Conductor (li2025image) AnimateDiff (256P) 2.98 11.30 0.214 0.664 91.64 10.29 0.192 0.644 31.22
Go-With-The-Flow burgert2025go CogVideoX-5B (480P) 0.60 15.62 0.392 0.490 41.99 14.59 0.410 0.425 10.27
Diffusion-As-Shader (gu2025diffusion) CogVideoX-5B (480P) 0.29 15.80 0.372 0.483 40.23 14.51 0.382 0.437 18.76
ATI (wang2025ati) Wan 2.1-14B (480P) 0.23 15.33 0.374 0.473 17.41 16.04 0.502 0.366 6.12
@@
Motion Transfer. We evaluate motion-following capability on two datasets: the DAVIS (Perazzi2016) validation set (30 videos) and 20 curated videos from the Sora webpage (videoworldsimulators2024). We use both because DAVIS presents challenging sequences with significant occlusions, while the Sora set provides clean examples with consistent visibility, ensuring a comprehensive evaluation. We directly compare the synthesized results with the corresponding ground-truth video frames. Visual fidelity is measured using PSNR, SSIM, and LPIPS (zhang2018unreasonable), while motion accuracy is assessed via End-Point Error (EPE), computed as the L2 distance between visible input tracks and the tracks extracted from the generated videos. All models are evaluated in its optimal configuration, and evaluations are performed at $832\times 480$ resolution after resizing. Speed measures are based on a single H100 GPU. We refer readers to Appendix for detailed protocols.
Camera Control. To assess our models’ ability to generate videos following camera controls, we evaluate its zero-shot performance on single-image 3D novel view synthesis. We compare it against several recent diffusion- and feed-forward-based view synthesis baselines (zhou2025stable; yu2024viewcrafter; xu2025depthsplat) on the LLFF dataset (mildenhall2019local). To adapt our 2D track-controlled model for this task, we first estimate scene geometry using a monocular depth network (wang2025moge) and compute a single scale factor to align the predicted depth with the reconstructed scene point cloud from COLMAP (schonberger2016structure). We then derive 2D motion trajectories by interpolating between the input and target frame using depth and camera parameters.
@@
DepthSplat (xu2025depthsplat) 576P 1.40 13.9 0.28 0.30
ViewCrafter (yu2024viewcrafter) 576P 0.26 14.0 0.30 0.30
SEVA (yu2024viewcrafter) 576P 0.20 14.1 0.30 0.29
@@
For all methods, the input image aspect ratio is fixed at 16:9, and all synthesized frames are resized to $512\times 288$ before evaluation with PSNR, SSIM, and LPIPS metrics. As shown in Table 2, our track-conditioned video generation model outperforms other 3D novel view synthesis baselines by a large margin, even though it is not specifically designed for this task. Moreover, our causal models achieve significantly higher generation throughput compared with both the baselines and their bidirectional counterparts.
@@ 4.3 Ablation Experiments
Track Representation. We compare our sinusoidal position encoding with a learnable track head against the RGB encoding strategy using a frozen VAE, following prior work (gu2025diffusion), where each 2D track is assigned a unique RGB color vector and placed onto a canvas before being fed into the VAE. Table 3 shows that our method (PE-Head) outperforms RGB-VAE in both efficiency and quality. Specifically, our lightweight PE-based encoding achieves better motion alignment while being two orders of magnitude faster than the VAE-based approach. We hypothesize sinusoidal encoding preserves stronger identification signals compared to RGB encoding due to richer expressive dimensions.
@@
We additionally show reconstruction accuracy and throughput as well as their time evolution in Table 4 and Figure 6(b,c) respectively. We denote our configuration as chunk-3, sink-1, window-1 (abbreviated c3s1w1), while sliding window approach used in Self Forcing is c3s0w6, attending to a maximum of 6 previous chunks without sink tokens, with unbounded RoPE positions that scale with temporal frame count. Although removing sink tokens (c3s0w1) yields marginal speed improvements, this modification comes at the cost of degraded long-term generation stability.
@@ Appendix A Training Efficient Tiny VAE
Tiny VAE (Wan 2.1  (BoerBohan2025TAEHV)) 9.84M 8×8×4 0.12 28.85 0.899 0.168
@@
Inspired by community implementations of Tiny VAE (BoerBohan2025TAEHV), we train a compact decoder from scratch with larger data, cleaner training pipelines, and better loss designs. While BoerBohan2025TAEHV trains Tiny VAE by regressing outputs to the original Wan VAE using adversarial loss from a PatchGAN discriminator and reconstruction loss with replay buffer, we extend this design by incorporating LPIPS loss with proper data scaling, and hyperparameter selections. We use a random subset of videos from OpenVid-1M and synthetic Wan videos (total 280K samples), training for 200K steps with a learning rate of $3\times 10^{-4}$, batch size of 16, and AdamW optimizer. We train at a lower resolution of $144\times 144$ and frame length of 21, but the model scales well to larger video dimensions.
@@ Appendix B VBench results and User study
We additionally evaluate MotionStream using VBench-I2V (huang2024vbench) and conduct user studies on 20 samples from the Sora demo subset. In Vbench-I2V, we exclude camera motion and dynamic degree metrics since these dimensions are already constrained by the input trajectory conditions rather than text prompts. VBench results show strong correlation with the underlying backbone model, favoring recent Wan-based architectures. The provision of image and trajectory conditions leads to uniformly high scores across methods, reducing discriminative power between models. Nonetheless, both our teacher and distilled models consistently achieve competitive performance across all evaluated dimensions.
@@
Image Conductor (li2025image) 0.847 0.868 0.791 0.889 0.906 0.505 0.689
GWTF (burgert2025go) 0.957 0.974 0.933 0.944 0.981 0.620 0.675
DAS (gu2025diffusion) 0.972 0.987 0.953 0.958 0.988 0.634 0.695
ATI (wang2025ati) 0.981 0.988 0.948 0.947 0.980 0.629 0.707
@@ Appendix C Additional Ablation Experiments and Qualitative Results
Impact of Chunk Size and Sampling Steps. In the main paper, we primarily focused on evaluating the efficiency of diverse sparse attention patterns with attention sink size and attention window size. As shown in Table 4 and Figure 6, reducing the chunk size increases the number of autoregressive rollouts required to generate the same length of video while reducing the amount of bidirectional attention, leading to worse performance. Since chunk sizes beyond 3 induce high latency, which makes them less useful for our interactive use case, we report only speed metrics for these configurations. While we did not perform controlled training ablations for larger chunk sizes, we observe similar trends with TalkingMachines (low2025talkingmachines), where quality generally improves with larger chunks at the cost of latency. Figure A2 (a) visualizes the latency and throughput given different chunk sizes and sampling steps. Note that since all benchmarks are conducted with a sink size and a window size of 1, the model with chunk size 7 has lower FPS than the one with chunk size 3 since its attention sequence length is longer.
@@
Table A4: Impact of Motion Control on Generative Quality. We evaluate whether injecting motion control degrades the pretrained model’s capability by comparing against a larger, dedicated I2V baseline (Wan 2.1 14B I2V). We also report performance when motion conditions are dropped. Results indicate that adding motion conditioning does not significantly degrade the base model’s generative quality. While removing motion conditions introduces a slight quality drop as our models were not optimized for this setting, the output still adheres to the given text and image inputs.
Model
i2v
subject
i2v
bg
subj.
consis.
bg
consis.
motion
smooth.
aesth.
qual.
imag.
qual.
FVD ($\downarrow$)
Wan 2.1 I2V (14B) 0.979 0.987 0.947 0.953 0.988 0.619 0.711 1274.6
Ours Teacher (1.3B) w. Motion 0.984 0.988 0.948 0.943 0.987 0.625 0.698 578.2
Ours Teacher (1.3B) w/o Motion 0.965 0.976 0.889 0.926 0.990 0.613 0.707 1550.4
Ours Student (1.3B) w. Motion 0.982 0.987 0.940 0.941 0.985 0.618 0.684 745.8
Ours Student (1.3B) w/o Motion 0.973 0.982 0.923 0.943 0.986 0.597 0.688 1532.5
Impact of Motion Control on Generative Capability. To assess whether injecting motion control compromises the base model’s generative capability, we compare our 1.3B models against the larger Wan 2.1 14B I2V baseline, which serves as a high-quality upper bound (as no Wan 1.3B I2V exists) on the Sora dataset. As shown in Table A4, we observe no noticeable degradation; despite the capacity gap and open-source training data, performance remains robust. In fact, a few demo samples demonstrate that motion control enables the model to generate dynamic motions that naive I2V models often struggle to produce due to static bias (choi2025enhancing). Consequently, FVD scores are lower with motion conditions, as the generated videos better mimic the target distribution through explicit motion constraints.
We also investigate the model’s behavior when motion conditions are not provided. Since our models are incentivized to accurately follow motion and are not explicitly trained for an I2V (or “motionless”) setup, providing empty motion inputs indeed results in a quality degradation. Nonetheless, while the quality is slightly lower compared to the fully conditioned setting, the model still follows text prompts effectively without severe visual collapse. One interesting observation is that the teacher model without motion conditions rarely produces sudden scene changes, yet results in a lower subject consistency metric. We hypothesize that applying CFG with empty motion conditions leads to unstable outputs where text prompts dominate. We did not observe this behavior with guidance-distilled student models. Please refer to the supplementary videos for detailed results.
@@
Streaming Demo. We show some examples of our streaming demo in Figure A5. The demo starts by accepting an input image and text prompts, which can help generate effects that are not achievable through mouse drags. Users can then choose the specific spacing/size of the track grids and start controlling objects in a scene, or move the camera. Due to its autoregressive nature, users can pause or resume the streaming generation process. Users can start/end or pause/resume the generation process with enter and space keyboard input, which is especially useful for dynamically adding static grids to specify unmoving regions or multiple moving grids to control different motions during streaming. As a video generative model, our method naturally supports drag-based image editing, generating intermediate transition frames as a bonus, while being faster than most dedicated drag-based image editing methods (pan2023drag; shi2024dragdiffusion; nie2023blessing; shin2024instantdrag; zhao2024fastdrag). To further support diverse downstream tasks, we will continuously update our front-end UI with additional features.
@@ Appendix D Training Details
Data Preprocessing. We train our models on two primary data sources: OpenVid-1M (nan2024openvid) and synthetic videos generated by larger Wan text-to-video models. For OpenVid-1M, we filter the dataset to 0.6M videos by requiring a minimum of 81 frames and a 16:9 aspect ratio, sampling at 16 FPS. For synthetic data, we use 70K samples for Wan 2.1 (81 frames, 480P resolution, generated by Wan 2.1 14B using text prompts from VidProm (wang2024vidprom)) and 30K publicly available samples for Wan 2.2 (121 frames, 720P resolution, generated by Wan 2.2 5B from the FastVideo team (fastvideo2024)).
We extract motion trajectories from all videos using CoTracker3 (karaev2024cotracker3), tracking points on a $\times50$ uniform grid. Our training follows a two-stage design: initial training on OpenVid-1M to establish general motion conditioning capability, followed by fine-tuning on cleaner synthetic data to improve trajectory adherence and reduce artifacts from noisy real-world videos. For the distillation phase, we use only synthetic samples since the distillation process converges quickly and requires less data.
@@
with timestep shifting (wang2025wan), defined as $t’(k,t)=\frac{kt/1000}{1+(k-1)(t/1000)}$, where $k=6.0$. The model is trained to predict the expected velocity fields conditioned on noisy latents, denoising time steps, text prompts $c_{t}$ and motion control embedding $c_{m}$, with timestep-dependent weighting $w_{t}$:
@@
Following CausVid (yin2025causvid), we adapt to few-step causal inference through ODE trajectory regression. We generate 4,000 ODE trajectories from the teacher model and train for 2,000 steps. While the transition from bidirectional to causal attention requires significant adaptation, we find variations within causal patterns (e.g., different window sizes) are similar enough to learn jointly with single model. We therefore train with diverse sparse causal attention masks, creating a unified initialization that supports flexible self-rollout configurations. We maintain a batch size of 128 with a learning rate $2\times 10^{-6}$.
Self Forcing-Style Distillation with DMD.
Self Forcing distillation converges quickly at around 400 steps with a batch size of 64. We set learning rates to $2\times 10^{-6}$ for the generator and $4\times 10^{-7}$ for the critic (fake score function), with a 1:5 update ratio and gradient truncation as described in Sec. 3.
We employ AdamW optimizer (kingma2014adam; loshchilov2017decoupled) with mixed precision (bfloat16) and PyTorch’s FSDP. Exponential moving average (EMA) is applied during teacher training and DMD distillation. Wan 2.1 variants train at $832\times 480$ resolution while Wan 2.2 trains at $1280\times 704$. With 32 A100 GPUs, training the Wan 2.1 teacher model takes roughly 3 days with causal adaptation and distillation completing in 20 hours, while Wan 2.2 requires slightly longer training times.